10 (b) Find the mean and variance of binomial distribution.


1
Expert's answer
2021-03-31T16:15:44-0400

Solution:

Let X be the random variable following binomial distribution with parameters n and p.

"X\sim B(n,p)"

First, we find the mean or expectation of X.

From the definition of expectation:

"\mathrm{E}(X)=\sum x \operatorname{Pr}(X=x)"

Thus:

"\mathrm{E}(X)=\sum_{k=0}^{n} k\left(\begin{array}{l}

n \\

k

\end{array}\right) p^{k} q^{n-k}" [Definition of Binomial Distribution, with "p+q=1" ]"=\sum_{k=1}^{n} k\left(\begin{array}{l}

n \\

k

\end{array}\right)" "p^{k} q^{n-k} \quad [\because k=0, k\left(\begin{array}{l}n \\
k
\end{array}\right) p^{k} q^{n-k}=0]"

"=\sum_{k=1}^{n} n\left(\begin{array}{l}
n-1 \\
k-1
\end{array}\right) p^{k} q^{n-k}"

"[\text{Factors of Binomial Coefficient}: k\left(\begin{array}{l}n \\ k\end{array}\right)=n\left(\begin{array}{l}n-1 \\ k-1\end{array}\right)]"

"\begin{array}{l}
=n p \sum_{k=1}^{n}\left(\begin{array}{l}
n-1 \\
k-1
\end{array}\right) p^{k-1} q^{(n-1)-(k-1)} \quad \text { taking out } n p \text { and using }(n-1) \\
=n p \sum_{j=0}^{m}\left(\begin{array}{c}
m \\
j
\end{array}\right) p^{j} q^{m-j} \quad \text { putting } m=n-1, j=k-1
\end{array}"

"=n p" [Binomial Theorem and "p+q=1"]

Now, we find variance.

"\mathrm{E}(X^2)=\sum x^2 \operatorname{Pr}(X=x)"

"\begin{aligned}
\mathrm{E}\left(X^{2}\right) &=\sum_{k \geq 0}^{n} k^{2}\left(\begin{array}{l}
n \\
k
\end{array}\right) p^{k} q^{n-k} \\
&=\sum_{k=0}^{n} k n\left(\begin{array}{c}
n-1 \\
k-1
\end{array}\right) p^{k} q^{n-k} \\
&=n p \sum_{k=1}^{n} k\left(\begin{array}{c}
n-1 \\
k-1
\end{array}\right) p^{k-1} q^{(n-1)-(k-1)} \\
&=n p \sum_{j=0}^{m}(j+1)\left(\begin{array}{c}
m \\
j
\end{array}\right) p^{j} q^{m-j} \\
&=n p\left(\sum_{j=0}^{m} j\left(\begin{array}{c}
m \\
j
\end{array}\right) p^{j} q^{m-j}+\sum_{j=0}^{m}\left(\begin{array}{c}
m \\
j
\end{array}\right) p^{j} q^{m-j}\right) \\
&=n p\left(\sum_{j=0}^{m} m\left(\begin{array}{c}
m-1 \\
j-1
\end{array}\right) p^{j} q^{m-j}+\sum_{j=0}^{m}\left(\begin{array}{c}
m \\
j
\end{array}\right) p^{j} q^{m-j}\right) \\
&=n p\left((n-1) p \sum_{j=1}^{m}\left(\begin{array}{c}
m-1 \\
j-1
\end{array}\right) p^{j-1} q^{(m-1)-(j-1)}+\sum_{j=0}^{m}\left(\begin{array}{c}
m \\
j
\end{array}\right) p^{j} q^{m-j}\right) \\
&=n p((n-1) p+1) \\
&=n^{2} p^{2}+n p(1-p)
\end{aligned}"

Now, "var(X)=E(X^2)-(E(X))^2"

"=n^2p^2+np(1-p)-(np)^2
\\ =n^2p^2+np(1-p)-n^2p^2
\\=np(1-p)"

Hence, mean = np, variance = np(1-p).